OpenStudy (anonymous):

Can anyone help me understand this sum? sum from k=0 to 6 of k(6 choose k)(1/6)^k(5/6)^(6-k). My professor said that we should try to solve it without simply calculating each and every term.

OpenStudy (anonymous):

\[\sum_{k=0}^{6}k \left(\begin{matrix}6 \\ k\end{matrix}\right)(\frac{ 1 }{ 6 })^k(\frac{ 5 }{ 6 })^{(6-k)}\]

OpenStudy (zarkon):

\[\sum_{k=0}^{n}k{n\choose k}\left(p\right)^{k}\left(1-p\right)^{n-k}\] \[=\sum_{k=1}^{n}k{n\choose k}\left(p\right)^{k}\left(1-p\right)^{n-k}\] \[=\sum_{k=1}^{n}k\frac{n!}{k!(n-k)!}\left(p\right)^{k}\left(1-p\right)^{n-k}\] \[=\sum_{k=1}^{n}\frac{n!}{(k-1)!(n-k)!}\left(p\right)^{k}\left(1-p\right)^{n-k}\] \[=\sum_{k=1}^{n}\frac{n(n-1)!}{(k-1)!(n-k)!}\left(p\right)^{k}\left(1-p\right)^{n-k}\] \[=\sum_{k=1}^{n}\frac{n(n-1)!}{(k-1)!(n-k)!}p\left(p\right)^{k-1}\left(1-p\right)^{n-k}\] \[=np\sum_{k=1}^{n}\frac{(n-1)!}{(k-1)!(n-k)!}\left(p\right)^{k-1}\left(1-p\right)^{n-k}\] let \(m=k-1\) \[=np\sum_{m=0}^{n-1}\frac{(n-1)!}{m!(n-k)!}\left(p\right)^{m}\left(1-p\right)^{n-k}\] \[=np\sum_{m=0}^{n-1}\frac{(n-1)!}{m!(n-1-(k-1))!}\left(p\right)^{m}\left(1-p\right)^{n-1-(k-1)}\] \[=np\sum_{m=0}^{n-1}\frac{(n-1)!}{m!(n-1-m)!}\left(p\right)^{m}\left(1-p\right)^{n-1-m}\] \[=np\sum_{m=0}^{n-1}{n-1\choose m}\left(p\right)^{m}\left(1-p\right)^{n-1-m}\] \[=np\cdot 1=np\] for you \(n=6\) and \(p=1/6\) thus \(np=1\)